OpenStudy (anonymous):

The radii of two circles are in the ratio of 3 to 1. Find the area of the smaller circle if the area of the larger circle is 27 sq. in. i think the answer is 6 pi correct me if im wrong

OpenStudy (anonymous):

\[\frac{r_2}{r_1}=3\] \[\frac{A_2}{A_1}=\frac{r_2}{r_1}=3^2\] \[A_2=9A_1\implies A_1=\frac{A_2}{9}=\frac{27}{9}\]

OpenStudy (anonymous):

so it'd be 9 pi?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

yeah @completeidiot

OpenStudy (anonymous):

hmmm... 3 pi..?

OpenStudy (anonymous):

the answer will not have pi in the solution the area of a circle is \[A=\pi r^2\] Jonask solved the problem by using ratios so if we take the ratio of the larger area to the smaller area, which we will denote as A1 and A2 \[\frac{A_1}{A_2} = \frac{\pi r_1^2}{\pi r_2 ^2 }\] then the pi cancels out \[\frac{A_1}{A_2} = (\frac{ r_1}{r_2 })^2\]

OpenStudy (anonymous):

since the proportion of smaller area to larger area is 1 to 9 that means the smaller area is 1/9 the area of the larger area and since the larger area is 27 in order to solve for the smaller area, you would need to divide 27 by 9 which is also shown by Jonask

OpenStudy (anonymous):

pi will not be part of the answer simply due to the fact that it was cancelled out when determining the ratio